consider the law of conservation of mass as you answer this problem. when ammonium nitrate breaks down explosively, it forms nitrogen gas, oxygen gas, and water. when 40 grams of ammonium nitrate explode, 14 grams of nitrogen and 8 grams of oxygen are formed. how many grams of water are formed?
1
Expert's answer
2011-10-20T08:17:25-0400
Mass of source – ammonium nitrate, is equal to sum of masses of its derivatives –
oxygen, nitrogen and water. We can write:
m(NH4NO3) = m(H2O) + m(N2) + m(O2)
m(H2O) = m(NH4NO3) - m(N2) - m(O2) = 40-14-8 = 18 g.